Two rockets are to be launched in the same direction two hours apart . the first rocket is launched at 9:00am and travels at 15000mph. If the second rocket travels at 22500 mph, at what time will the two rockets be at the same distance from earth?
Answer by Alan3354(69443)

Two rockets are to be launched in the same direction two hours apart . the first rocket is launched at 9:00am and travels at 15000mph. If the second rocket travels at 22500 mph, at what time will the two rockets be at the same distance from earth?
--------------
The 2nd rocket is 22500 - 15000 = 7500 mi/hr faster.
The 1st rocket is 15000 mi/hr * 2 hr = 30000 miles ahead.
------
30000/7500 = 4 hours.
----
0900 + 2 + 4 = 1500 for the same distance.
or 3 PM.